OpenStudy (anonymous):

If B=17 degrees 55', C=98 degrees 15' and a=17, find c?

OpenStudy (campbell_st):

you can find angle A using angle sum of a triangle. Then you need the Law of sine to find c \[\frac{a}{\sin(A)} = \frac{c}{\sin(C)}\] hope it helps

OpenStudy (triciaal):

 then as above for c